OpenStudy (anonymous):

a tank contain 1000 liters water. Brine that contain 0.05kg of salt per liter of water enters the tank at a rate of 8 liter per minute. Brine that contain 0.04kg of salt per liter of water enters the tank at a rate of 5 liter per minute. The well-mixed solution drains from the tank at a rate of 13 liters per minute. 1. write differential equation that the amount of salt, y(t) must satisfy. 2. find the amount of salt in the tank at any time, t. 3. how much salt is present after 20 minutes? 4. what is the concentration of solution in the tank after a very long time?

OpenStudy (paxpolaris):

1. \(dy/dt=0.05∗8+0.04∗5−13y/1000\)

OpenStudy (anonymous):

thanks, can you also solve the rest?

OpenStudy (paxpolaris):

no, i'm probably wrong

OpenStudy (experimentx):

Ah ... y is salt right??

OpenStudy (paxpolaris):

mass of salt in tank

OpenStudy (experimentx):

i think it is right ///

OpenStudy (experimentx):

these two are inputs 0.05∗8+0.04∗5 and this y13/1000 is decay constant

OpenStudy (experimentx):

13/1000 is decay constant

OpenStudy (experimentx):

I'm pretty sure this must be decay equation ... that must converge to input

OpenStudy (experimentx):

to find the amount of salt at any time t, solve equation 1) y(t) = some function of t + some constant ----- (2 you will need an initial condition, assuming that there is no salt present initially, y(0) = 0 <--- using this find the value of 'some constant' then put the value of 'some constant' ... in above ... which is your answer of Q.2 Q.3 put t=20 in y(20) = ...

OpenStudy (experimentx):

I am guessing that the Q.4 should be 0.046

OpenStudy (paxpolaris):

yes

OpenStudy (experimentx):

Oh ... gotta sleep, good night!

OpenStudy (anonymous):

Could you please show me all the formula?

OpenStudy (paxpolaris):

simplifying differential equation for Q1.: \[\large y\ '(t )= 0.6-0.013y\]

OpenStudy (paxpolaris):

\[\large {dy \over dt} = {0.6 -0.013y }\]\[\implies \large {dy \over 0.6-0.013y} = dt\] Integrate both sides, and get rid of the constant using \(\large y(0)=0\)

OpenStudy (paxpolaris):

that will give you the equation for Q2

OpenStudy (anonymous):

ln(0.6-0.013y)=t+c. ?

OpenStudy (paxpolaris):

\[\Large -\frac 1 {0.013}\ln(0.6-0.013y)=t+C_1\]

OpenStudy (anonymous):

Q3=10.5668kg?

OpenStudy (anonymous):

How about Q4? Not 46?

OpenStudy (paxpolaris):

46.15 kg in 1000 liters

OpenStudy (paxpolaris):

so, concentration will approach 0.04615 kg/l

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Can you solve my another closed problem?

OpenStudy (paxpolaris):

post link

OpenStudy (anonymous):

dy/dx=y^3-2y^2-15y What are the constant solutions? For what values of y is y(x) increasing? And what decreasing? Use the information groom above to sketch the direction field for the given differential equation